OpenStudy (anonymous):

integrat 36/4x^2 -12x + 9

OpenStudy (anonymous):

maybe we can get lucky and factor the denominator

OpenStudy (abb0t):

you can factor the denominator. common number being -6.

OpenStudy (anonymous):

try \[\int \frac{dx}{(2x-3)^2}\]

OpenStudy (abb0t):

to make it eaiser you can factor, or you can leave as is :)

OpenStudy (anonymous):

then use \(u=2x-3\) and you should be done quickly

sam (.sam.):

and 36 satellite

OpenStudy (anonymous):

36 is unimportant in the computation, just remember to stick it in at the end

sam (.sam.):

yeah

OpenStudy (anonymous):

why 36 is unimportant?

sam (.sam.):

Because you are sticking it out \[36[\int\limits \frac{dx}{(2x-3)^2}]\]

OpenStudy (anonymous):

OH OK THANKYOU v. much